OpenStudy (anonymous):

Why does the inequality sign have to be flipped when multiplying and dividing by a negative number like in -2x>10? PLZ HELP

OpenStudy (anonymous):

help? @Loser66

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I think you flip the sign because the value changes to negative.....I just do the rule of flipping it. Never really thought about the explanation

OpenStudy (anonymous):

here is one explanation it relies on defining \(a<b\) and there is a positive \(p\) with \(a+p=b\)

OpenStudy (anonymous):

for example \(-2<1\) because \(-2+3=1\)

OpenStudy (anonymous):

now multiply by a negative number call it \(n\) \[n(a+p)=nb\\ na+np =nb\]

OpenStudy (anonymous):

now since \(p\) is positive and \(n\) is negative that means \(np\) is negative

OpenStudy (anonymous):

im sorry im just confused

OpenStudy (loser66):

:( me too.

OpenStudy (anonymous):

so the inequality has been switched as now \[na=nb-np\] and \(np>0\)

OpenStudy (anonymous):

if you want a satisfactory explanation of why you change the sense of the inequality, you need an actual definition for what \(a<b\) means

OpenStudy (anonymous):

an example may make it clear we say \(-2<1\) because \(-2+3=1\) and \(3>0\)

OpenStudy (anonymous):

-2x>10 this is hte example i was given in the question

OpenStudy (anonymous):

now multiply \[-2+3=1\] on both sides by a number, i pick \(-5\)

OpenStudy (anonymous):

you get \[10-15=-5\] or \[10=-5+15\] making \(10>-5\) and the inequality has been switched

OpenStudy (anonymous):

do you get it? @Loser66

OpenStudy (loser66):

How to say!! Have to confess that I got until a<b meaning.

OpenStudy (anonymous):

i dont even know at this point what to put lol

OpenStudy (loser66):

ok, I got what he meant now. Wow!!

OpenStudy (loser66):

hahaha... it's fun when I don't know how to explain but I have to explain the explanation of other.

OpenStudy (anonymous):

haha

OpenStudy (loser66):

@arianaharmony Your question is about "Why we have to change the sign when multiple/ divide both sides of an inequality with a negative number?" right? It's not "solve -2x>10" right? So that, let the original problem aside. just answer your question, ok??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

It's only flipped when you divide by a negative

OpenStudy (loser66):

he gave us an example -2<1 and definition of it, it says " minus two is called lesser than one when there is a number such that minus two plus that number equal to one" got this part?

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

oh ok i think i might get it

OpenStudy (loser66):

In real life: Loser is poor, all his property is 5 dollars arianaharmony is rich, he has 8 dollars so that Loser's money is < arianaharmony 's money. Why I can say so?? Yes I can say that because Loser's money + some more = arianaharmony's money Proof 5 + 3 =8

OpenStudy (loser66):

so that "some more" is 3, got me?

OpenStudy (anonymous):

ya ok i get that

OpenStudy (loser66):

hahaha... let big one do big thing, small one do small thing, I am happy to be small one to be here to explain this stuff to you. lol

OpenStudy (loser66):

now, back to his example, so -2<1 when -2+ "some more" =1, that "some more" is 3 , OK? -2+3=1

OpenStudy (anonymous):

yes okay. but after this question would you mind helping me with 2 other questions? ik its alot but i need help lol sorry

OpenStudy (loser66):

Actually, if you can understand the stuff, it's perfect. but if you don't, memorize the rule is perfect, too. hihihi... Now, what are your problems?

OpenStudy (anonymous):

wait. what was the answer to this question exactly?

OpenStudy (loser66):

WAAAT?? ok -2x>10 divided both sides by -2 and change the sign, you have x<5

OpenStudy (anonymous):

but the question is Why does the inequality sign have to be flipped when multiplying and dividing by a negative number

OpenStudy (anonymous):

if you dont know its okay. can you help me with my other questions tho plz @Loser66

OpenStudy (loser66):

I give you the link to get a neat answer, hahaha.... click at the question you want to know the answer. The Answerer is a Ph.D http://mathforum.org/library/drmath/sets/select/dm_inequal_neg.html

OpenStudy (loser66):

lalalala.....lallala. life is beautiful. hahaha

OpenStudy (anonymous):

can you help me with this problem? Explain the steps in solving the following absolute value equation. 2|x-2|=12.

OpenStudy (loser66):

close the post, open a new one to get help. I told you, I don't know how to explain. I am sorry for that.

OpenStudy (anonymous):

ok thanks